OpenStudy (anonymous):

find the exact solution to the equation 64(1/4)^x/2=4 help?

OpenStudy (anonymous):

64 (1/4)^(x/2)=4

OpenStudy (anonymous):

(1/4)^(x/2)=(4/64)

OpenStudy (anonymous):

use log...

OpenStudy (anonymous):

log(1/4)^(x/2)=log(4/64) (x/2)log(1/4)=log(4/64) x/2=log(4/64)/log(1/4)=log(4/64-1/4) x=2*log(4/64-1/4) x=2*log((4-16)/64) x=log(12/64)^2